OpenStudy (anonymous):

divide: 36xy-397/3y explain

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

so i change the negative sign to a positive?

OpenStudy (danjs):

k

OpenStudy (anonymous):

wait how do i clt if you have to divide 3?

OpenStudy (danjs):

are both terms divided by 3y or just the second?

OpenStudy (anonymous):

well whatever the sum for the numerator is, is supposed to be divided by 3

OpenStudy (jhannybean):

divide each term in the numerator by \(3y\)

OpenStudy (anonymous):

so do i go -362/3???

OpenStudy (danjs):

\[\frac{ 36xy - 397 }{ 3y } = \frac{ 36 }{ 3 }*\frac{ xy }{ y } - \frac{ 397 }{ 3 }*\frac{ 1 }{ y }\]

OpenStudy (jhannybean):

\[\frac{36xy-397}{3y}\]That is the question, yes?

OpenStudy (danjs):

divide the numbers and the variables separately like above

OpenStudy (anonymous):

well 36-397 = -361

OpenStudy (jhannybean):

You cannot do that.

OpenStudy (jhannybean):

Big nono.

OpenStudy (danjs):

36/3 = 6 (xy)/y = x 397/3 = 397/3

OpenStudy (anonymous):

ohh okay i divide each by 3....

OpenStudy (jhannybean):

yes,

OpenStudy (danjs):

Remember before, you can only add terms that have the same powers on the x

OpenStudy (anonymous):

36/3=12 397/3=132.3

OpenStudy (jhannybean):

Keep the fractions that do not simply as fractions instead of decimals unless otherwise asked

OpenStudy (jhannybean):

simplify*

OpenStudy (anonymous):

im confused.....

OpenStudy (jhannybean):

So leave \(\dfrac{397}{3}\) alone.

OpenStudy (anonymous):

okay so 36/3=12 still

OpenStudy (jhannybean):

Yes.

OpenStudy (danjs):

\[\frac{ 36xy }{ 3y } = \frac{ 36 }{ 3 }*\frac{ xy }{ y } = 3 * \frac{ x*y*y ^{-1} }{ 1 } = 3 * x*y ^{1-1}\]

OpenStudy (danjs):

3x for that one

OpenStudy (danjs):

12x... lol i cant divide

OpenStudy (jhannybean):

:P

OpenStudy (anonymous):

omg i forgot about the 1-1 lol

OpenStudy (danjs):

yes if you are dividing variables to pwers, just remember if you move one to the other side of the fraction, subtract its power

OpenStudy (danjs):

\[\frac{ y^6 }{ y^3 } = y ^{6-3} = y^3\] for example

OpenStudy (jhannybean):

I would go about it this way:\[\frac{36xy-397}{3y}\]\[\left(\frac{\color{blue}{36}x\color{red}y}{\color{blue}{3}\color{red}{y}}\right)-\frac{397}{3y}\]

OpenStudy (jhannybean):

So simplify the blue, simplify the red, and you're done.

OpenStudy (danjs):

im goin to play skyrim.... later

OpenStudy (jhannybean):

xD

OpenStudy (jhannybean):

Whichever method makes sense to you, stick to it and solve it.

OpenStudy (jhannybean):

So for my method (36/3) = 12, right? (y/y) = 1, yes?

OpenStudy (jhannybean):

\[\left(\frac{\color{blue}{36}x\color{red}y}{\color{blue}{3}\color{red}{y}}\right)-\frac{397}{3y} = 12x -\frac{397}{3y}\]